# Cross-correlation between the curvature perturbations and magnetic fields in pure ultra slow roll inflation

### v1

@@ Appendix A Does the energy density of the electromagnetic field remain positive?
Earlier, in the case of the action (34) containing the non-conformal coupling of the form $J^{2}(X)\simeq J_{0}\,X^{n/2}$, we had obtained the corresponding energy density $\rho_{{}_{\mathrm{EM}}}$ of the electromagnetic field to be given by Eq. (39). While the expectation values $\langle\hat{\rho}_{{}_{\mathrm{B}}}\rangle$ and $\langle\hat{\rho}_{{}_{\mathrm{E}}}\rangle$ are clearly positive definite [cf. Eqs. (22), (23) and (24)], the quantity $\langle\hat{\rho}_{{}_{\mathrm{{EM}}}}\rangle$ contains $n$ which may have either sign. This may raise the concern as to whether the total energy density of the electromagnetic field remains positive definite. To address the concern, let us study the behavior of the energy density of the electromagnetic field over the entire span of conformal time of our interest. For our choice of the non-conformal coupling function, i.e. $J=(\eta/\eta_{\mathrm{e}})^{n\,p/2}$ [cf. Eqs. (37) and (6)], we obtain the quantity ${\mathcal{A}}_{k}$ describing the Fourier modes of the electromagnetic vector potential to be
@@ B.1 Case of $p=6$
${\mathcal{G}}_{2}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=-\frac{iH_{{}_{\mathrm{I}}}}{256\,M_{{}_{\mathrm{Pl}}}\,\sqrt{k1^{3}\,k_{2}^{5}\,k_{3}^{5}\,\epsilon_{1\mathrm{i}}}\,\eta_{\mathrm{e}}^{3}}\,\left(\frac{\eta_{\mathrm{i}}}{\eta_{\mathrm{e}}}\right)^{3}\,\left[k_{1}^{4}-2\,k_{1}^{2}\,(k_{2}^{2}+k_{3}^{2})+k_{2}^{4}+6\,k_{2}^{2}\,k_{3}^{2}+k_{3}^{4}\right]$
($\quad\biggl{(}\biggl{\{}-24\,i-24\,(k_{1}+k_{2}+k_{3})\,\eta_{\mathrm{e}}-6\,i\,\left[k_{1}^{2}-4\,k_{1}\,(k_{2}+k_{3})-k_{2}^{2}-4\,k_{2}\,k_{3}-k_{3}^{2}\right]\,\eta_{\mathrm{e}}^{2}$
@@
×$\quad\times\mathrm{Ei}[i\,(k_{1}+k_{2}+k_{3})\,\eta_{\mathrm{e}}]\biggr{)},$ (83b)
${\mathcal{G}}_{3}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=-\frac{i\,H_{{}_{\mathrm{I}}}}{96\,M_{{}_{\mathrm{Pl}}}\,k_{1}\,\sqrt{k_{1}\,k_{2}\,k_{3}\,\epsilon_{1\mathrm{i}}}\,\eta_{\mathrm{e}}}\,\left(\frac{\eta_{\mathrm{i}}}{\eta_{\mathrm{e}}}\right)^{3}\,(k_{1}^{2}-k_{2}^{2}-k_{3}^{2})\,\biggl{(}\biggl{\{}6\,i+6\,(k_{1}+k_{2}+k_{3})\,\eta_{\mathrm{e}}$
$\quad-i\,\left[k_{1}^{4}+2\,k_{1}^{2}\,(k_{2}^{2}+k_{3}^{2})-3\,(k_{2}^{2}-k_{3}^{2})^{2}\right]\,\eta_{\mathrm{e}}^{4}\;\mathrm{Ei}(i(k_{1}+k_{2}+k_{3})\eta_{\mathrm{e}})\biggr{)},$ (83c)
${\mathcal{G}}_{4}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=\frac{7\,i\,H_{{}_{\mathrm{I}}}}{768\,M_{{}_{\mathrm{Pl}}}\,\sqrt{k_{1}^{3}\,k_{2}^{5}\,k_{3}^{5}\,\epsilon_{1\mathrm{i}}}\,\eta_{\mathrm{e}}^{3}}\left(\frac{\eta_{\mathrm{i}}}{\eta_{\mathrm{e}}}\right)^{3}\,\left[k_{1}^{4}-2\,k_{1}^{2}\,(k_{2}^{2}+k_{3}^{2})+k_{2}^{4}+6\,k_{2}^{2}\,k_{3}^{2}+k_{3}^{4}\right]$
@@
×$\quad\times\,\mathrm{Ei}[i\,(k_{1}+k_{2}+k_{3})\,\eta_{\mathrm{e}}]\biggr{)},$ (83d)
${\mathcal{G}}_{5}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=-\frac{i\,H_{{}_{\mathrm{I}}}}{128\,M_{{}_{\mathrm{Pl}}}\,\sqrt{k_{1}^{3}\,k_{2}\,k_{3}^{5}\,\epsilon_{1\mathrm{i}}}\,\eta_{\mathrm{e}}}\,\left(\frac{\eta_{\mathrm{i}}}{\eta_{\mathrm{e}}}\right)^{3}\,\left[2\,k_{3}^{2}\,(k_{1}^{2}+k_{2}^{2})+(k_{1}^{2}-k_{2}^{2})^{2}-3\,k_{3}^{4}\right]$
($\quad\biggl{(}\biggl{\{}-6\,i-6\,(k_{1}+k_{2}+k_{3})\,\eta_{\mathrm{e}}-i\,\left[-6\,k_{3}\,(k_{1}+k_{2})+3\,(k_{1}-k_{2})^{2}-k_{3}^{2}\right]\,\eta_{\mathrm{e}}^{2}$
$\quad+i\left[2\,k_{3}^{2}\,(k_{1}^{2}+k_{2}^{2})-3\,(k_{1}^{2}-k_{2}^{2})^{2}+k_{3}^{4}\right]\,\eta_{\mathrm{e}}^{4}\;\mathrm{Ei}[i\,(k_{1}+k_{2}+k_{3})\,\eta_{\mathrm{e}}]\biggr{)}+{\bm{k}}_{2}\leftrightarrow{\bm{k}}_{3},$ (83e)
${\mathcal{G}}_{6}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=-\frac{i\,H_{{}_{\mathrm{I}}}^{2}\,\sqrt{\epsilon_{1\mathrm{i}}}}{16\,M_{{}_{\mathrm{Pl}}}\,(k_{1}+k_{2}+k_{3})^{6}\,\sqrt{k_{1}^{7}\,k_{2}\,k_{3}^{5}}\,\eta_{\mathrm{i}}^{3}}\,\mathrm{e}^{i\,(k_{1}+k_{2}+k_{3})\,\eta_{\mathrm{e}}}\,\left[2\,k_{3}^{2}\,(k_{1}^{2}+k_{2}^{2})+(k_{1}^{2}-k_{2}^{2})^{2}-3\,k_{3}^{4}\right]$
×$\quad\times\biggl{\{}-6\,\biggl{[}4\,k_{1}^{5}+24\,k_{1}^{4}\,(k_{2}+k_{3})+8\,k_{1}^{3}\,(6\,k_{2}^{2}+15\,k_{2}k_{3}+7\,k_{3}^{2})$
$\quad+k_{1}^{2}\,(43\,k_{2}^{3}+168\,k_{2}^{2}\,k_{3}+216\,k_{2}\,k_{3}^{2}+56\,k_{3}^{3})$
$\quad+6\,k_{1}\,(k_{2}+k_{3})\,(3\,k_{2}^{3}+12\,k_{2}^{2}\,k_{3}+16\,k_{2}\,k_{3}^{2}+4\,k_{3}^{3})+(k_{2}+k_{3})^{2}\,(3\,k_{2}^{3}+12\,k_{2}^{2}\,k_{3}+16\,k_{2}\,k_{3}^{2}+4\,k_{3}^{3})\biggr{]}$
$\quad+3\,i\,\bigg{[}5\,k_{1}^{5}+33\,k_{1}^{4}\,(k_{2}+k_{3})+2\,k_{1}^{3}\,(33\,k_{2}^{2}+90\,k_{2}\,k_{3}+41\,k_{3}^{2})$
$\quad+k_{1}^{2}\,(56\,k_{2}^{3}+246\,k_{2}^{2}\,k_{3}+342\,k_{2}\,k_{3}^{2}+82\,k_{3}^{3})+3\,k_{1}\,(k_{2}+k_{3})\,(7\,k_{2}^{3}+33\,k_{2}^{2}\,k_{3}+49\,k_{2}\,k_{3}^{2}+11\,k_{3}^{3})$
$\quad+(k_{2}+k_{3})^{2}\,(3\,k_{2}^{3}+15\,k_{2}^{2}\,k_{3}+23\,k_{2}\,k_{3}^{2}+5\,k_{3}^{3})\bigg{]}\,(k_{1}+k_{2}+k_{3})\,\eta_{\mathrm{e}}$
$\quad+3\,\biggl{[}k_{1}^{5}+9\,k_{1}^{4}\,(k_{2}+k_{3})+2\,k_{1}^{3}\,(9\,k_{2}^{2}+30\,k_{2}\,k_{3}+13\,k_{3}^{2})+k_{1}^{2}\,(13\,k_{2}^{3}+78\,k_{2}^{2}k_{3}+126\,k_{2}\,k_{3}^{2}+26\,k_{3}^{3})$
$\quad+3\,k_{1}\,(k_{2}+k_{3})\,(k_{2}^{3}+9\,k_{2}^{2}\,k_{3}+17\,k_{2}\,k_{3}^{2}+3\,k_{3}^{3})+k_{3}\,(k_{2}+k_{3})^{2}\,(3\,k_{2}^{2}+7\,k_{2}\,k_{3}+k_{3}^{2})\biggr{]}$
×$\quad\times(k_{1}+k_{2}+k_{3})^{2}\,\eta_{\mathrm{e}}^{2}$
$\quad-i\,\biggl{[}3\,k_{1}^{4}\,(k_{2}+k_{3})+k_{1}^{3}\,(6\,k_{2}^{2}+33\,k_{2}\,k_{3}+13\,k_{3}^{2})+k_{1}^{2}\,(3\,k_{2}^{3}+39\,k_{2}^{2}\,k_{3}+84\,k_{2}\,k_{3}^{2}+13\,k_{3}^{3})$
$\quad+3\,k_{1}\,k_{3}\,(k_{2}+k_{3})\,(3\,k_{2}^{2}+10\,k_{2}\,k_{3}+k_{3}^{2})+3\,k_{2}\,k_{3}^{2}\,(k_{2}+k_{3})^{2}\biggr{]}\,(k_{1}+k_{2}+k_{3})^{3}\,\eta_{\mathrm{e}}^{3}$
$\quad-k_{1}\,k_{3}\,\left[k_{1}^{2}\,(3\,k_{2}+k_{3})+k_{1}\,(3\,k_{2}^{2}+12\,k_{2}\,k_{3}+k_{3}^{2})+3\,k_{2}\,k_{3}\,(k_{2}+k_{3})\right]\,(k_{1}+k_{2}+k_{3})^{4}\,\eta_{\mathrm{e}}^{4}$
$\quad+i\,k_{1}^{2}\,k_{2}\,k_{3}^{2}\,(k_{1}+k_{2}+k_{3})^{5}\,\eta_{\mathrm{e}}^{5}\biggr{\}}+{\bm{k}}_{2}\leftrightarrow{\bm{k}}_{3},$ (83f)
@@ B.2 Case of $p=4$
${\mathcal{G}}_{2}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=-\frac{i\,H_{{}_{\mathrm{I}}}}{128\,M_{{}_{\mathrm{Pl}}}\,\sqrt{k_{1}\,k_{2}^{5}\,k_{3}^{5}\,\epsilon_{1\mathrm{i}}}\,\eta_{\mathrm{e}}^{2}}\,\left(\frac{\eta_{\mathrm{i}}}{\eta_{\mathrm{e}}}\right)^{2}\,\left[k_{1}^{4}-2\,k_{1}^{2}\,(k_{2}^{2}+k_{3}^{2})+k_{2}^{4}+6\,k_{2}^{2}\,k_{3}^{2}+k_{3}^{4}\right]$
@@
$\quad+\left[2\,k_{3}^{2}\,(k_{2}^{2}-3k_{1}^{2})+3\,(k_{1}^{2}-k_{2}^{2})^{2}+3\,k_{3}^{4}\right]\,\eta_{\mathrm{e}}^{4}\;\mathrm{Ei}[i\,(k_{1}+k_{2}+k_{3})\,\eta_{\mathrm{e}}]\biggr{)},$ (84b)
${\mathcal{G}}_{3}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=-\frac{i\,H_{{}_{\mathrm{I}}}}{8\,M_{{}_{\mathrm{Pl}}}\,\sqrt{k_{1}\,k_{2}\,k_{3}\,\epsilon_{1\mathrm{i}}}}\,\left(\frac{\eta_{\mathrm{i}}}{\eta_{\mathrm{e}}}\right)^{2}\,(k_{1}^{2}-k_{2}^{2}-k_{3}^{2})\,\biggl{\{}\biggl{[}1-i\,(k_{1}+k_{2}+k_{3})\,\eta_{\mathrm{e}}$
$\quad-\frac{2\,k_{1}\,k_{2}\,k_{3}\,\eta_{\mathrm{e}}^{2}}{(k_{1}+k_{2}+k_{3})}\biggr{]}\,\mathrm{e}^{i\,(k_{1}+k_{2}+k_{3})\,\eta_{\mathrm{e}}}-(k_{1}^{2}+k_{2}^{2}+k_{3}^{2})\,\eta_{\mathrm{e}}^{2}\;\mathrm{Ei}[i\,(k_{1}+k_{2}+k_{3})\,\eta_{\mathrm{e}}]\biggr{\}},$ (84c)
${\mathcal{G}}_{4}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=\frac{3\,i\,H_{{}_{\mathrm{I}}}}{128\,M_{{}_{\mathrm{Pl}}}\,\sqrt{k_{1}\,k_{2}^{5}\,k_{3}^{5}\,\epsilon_{1\mathrm{i}}}\,\eta_{\mathrm{e}}^{2}}\,\left(\frac{\eta_{\mathrm{i}}}{\eta_{\mathrm{e}}}\right)^{2}\,\left[k_{1}^{4}-2\,k_{1}^{2}\,(k_{2}^{2}+k_{3}^{2})+k_{2}^{4}+6\,k_{2}^{2}\,k_{3}^{2}+k_{3}^{4}\right]$

### v2

@@ Appendix A Does the energy density of the electromagnetic field remain positive?
Earlier, in the case of the action in Eq. (34), which contains the non-conformal coupling of the form $J^{2}(X)\simeq J_{0}\,X^{n/2}$, we had obtained the corresponding energy density $\rho_{{}_{\mathrm{EM}}}$ of the electromagnetic field to be given by Eq. (39). While the expectation values $\langle\hat{\rho}_{{}_{\mathrm{B}}}\rangle$ and $\langle\hat{\rho}_{{}_{\mathrm{E}}}\rangle$ are clearly positive definite [cf. Eqs. (22), (23) and (24)], the quantity $\langle\hat{\rho}_{{}_{\mathrm{{EM}}}}\rangle$ contains $n$ which may have either sign. This may raise the concern as to whether the total energy density of the electromagnetic field remains positive definite. To address the concern, let us study the behavior of the energy density of the electromagnetic field over the entire span of conformal time of our interest. For our choice of the non-conformal coupling function, i.e. $J=(\eta/\eta_{\mathrm{e}})^{n\,p/2}$ [cf. Eqs. (37) and (6)], we obtain the quantity ${\mathcal{A}}_{k}$ describing the Fourier modes of the electromagnetic vector potential to be
@@ B.1 Case of $p=6$
${\mathcal{G}}_{2}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=-\frac{i\,H_{{}_{\mathrm{I}}}}{96\,M_{{}_{\mathrm{Pl}}}\,k_{1}\,\sqrt{k_{1}\,k_{2}\,k_{3}\,\epsilon_{1\mathrm{i}}}\,\eta_{\mathrm{e}}}\,\left(\frac{\eta_{\mathrm{i}}}{\eta_{\mathrm{e}}}\right)^{3}\,(k_{1}^{2}-k_{2}^{2}-k_{3}^{2})\,\biggl{(}\biggl{\{}6\,i+6\,(k_{1}+k_{2}+k_{3})\,\eta_{\mathrm{e}}$
$\quad-i\,\left[k_{1}^{4}+2\,k_{1}^{2}\,(k_{2}^{2}+k_{3}^{2})-3\,(k_{2}^{2}-k_{3}^{2})^{2}\right]\,\eta_{\mathrm{e}}^{4}\;\mathrm{Ei}(i(k_{1}+k_{2}+k_{3})\eta_{\mathrm{e}})\biggr{)},$ (83b)
${\mathcal{G}}_{3}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=-\frac{i\,H_{{}_{\mathrm{I}}}}{128\,M_{{}_{\mathrm{Pl}}}\,\sqrt{k_{1}^{3}\,k_{2}\,k_{3}^{5}\,\epsilon_{1\mathrm{i}}}\,\eta_{\mathrm{e}}}\,\left(\frac{\eta_{\mathrm{i}}}{\eta_{\mathrm{e}}}\right)^{3}\,\left[2\,k_{3}^{2}\,(k_{1}^{2}+k_{2}^{2})+(k_{1}^{2}-k_{2}^{2})^{2}-3\,k_{3}^{4}\right]$
×($\quad\times\biggl{(}\biggl{\{}-6\,i-6\,(k_{1}+k_{2}+k_{3})\,\eta_{\mathrm{e}}-i\,\left[-6\,k_{3}\,(k_{1}+k_{2})+3\,(k_{1}-k_{2})^{2}-k_{3}^{2}\right]\,\eta_{\mathrm{e}}^{2}$
$\quad+i\left[2\,k_{3}^{2}\,(k_{1}^{2}+k_{2}^{2})-3\,(k_{1}^{2}-k_{2}^{2})^{2}+k_{3}^{4}\right]\,\eta_{\mathrm{e}}^{4}\;\mathrm{Ei}[i\,(k_{1}+k_{2}+k_{3})\,\eta_{\mathrm{e}}]\biggr{)}+{\bm{k}}_{2}\leftrightarrow{\bm{k}}_{3},$ (83c)
${\mathcal{G}}_{4}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=-\frac{H_{{}_{\mathrm{I}}}\,\sqrt{\epsilon_{1\mathrm{i}}}}{16\,M_{{}_{\mathrm{Pl}}}\,\sqrt{k_{1}^{7}k_{2}k_{3}^{5}}\,(k_{1}+k_{2}+k_{3})^{5}\,\eta_{\mathrm{i}}^{3}}\,\left[(k_{1}^{2}-k_{2}^{2})^{2}+2\,(k_{1}^{2}+k_{2}^{2})\,k_{3}^{2}-3\,k_{3}^{4}\right]$
×($\quad\times\biggl{(}\biggl{\{}3\,\biggl{[}4\,k_{1}^{5}+23\,k_{1}^{4}(k_{2}+k_{3})+k_{1}^{3}\,\left(46\,k_{2}^{2}+107\,k_{2}\,k_{3}+51\,k_{3}^{2}\right)$
$\quad+k_{1}^{2}\,\left(42\,k_{2}^{3}+153\,k_{2}^{2}\,k_{3}+178\,k_{2}\,k_{3}^{2}+51\,k_{3}^{3}\right)$
$\quad+k_{1}\,(k_{2}+k_{3})\,\left(18\,k_{2}^{3}+69\,k_{2}^{2}\,k_{3}+84\,k_{2}\,k_{3}^{2}+23\,k_{3}^{3}\right)+(k_{2}+k_{3})^{2}\,\left(3\,k_{2}^{3}+12\,k_{2}^{2}\,k_{3}+15\,k_{2}\,k_{3}^{2}+4\,k_{3}^{3}\right)\biggr{]}$
$\quad-3\,i\,(k_{1}+k_{2}+k_{3})\,\biggl{[}k_{1}^{5}+8\,k_{1}^{4}\,(k_{2}+k_{3})+k_{1}^{3}\,\left(16\,k_{2}^{2}+47\,k_{2}\,k_{3}+21\,k_{3}^{2}\right)$
$\quad+k_{1}^{2}\,(k_{2}+3\,k_{3})\,\left(12\,k_{2}^{2}+27\,k_{2}\,k_{3}+7\,k_{3}^{2}\right)+k_{1}\,(k_{2}+k_{3})\,\left(3\,k_{2}^{3}+24\,k_{2}^{2}\,k_{3}+39\,k_{2}\,k_{3}^{2}+8\,k_{3}^{3}\right)$
$\quad+k_{3}\,(k_{2}+k_{3})^{2}\,\left(3\,k_{2}^{2}+6\,k_{2}\,k_{3}+k_{3}^{2}\right)\biggr{]}\,\eta_{\mathrm{e}}-3\,(k_{1}+k_{2}+k_{3})^{2}\,\biggl{[}k_{1}^{4}\,(k_{2}+k_{3})+2\,k_{1}^{3}\,\left(k_{2}^{2}+5\,k_{2}\,k_{3}+2\,k_{3}^{2}\right)$
$\quad+k_{1}^{2}\left(k_{2}^{3}+12\,k_{2}^{2}\,k_{3}+23\,k_{2}\,k_{3}^{2}+4\,k_{3}^{3}\right)+k_{1}\,k_{3}\,(k_{2}+k_{3})\,\left(3\,k_{2}^{2}+9\,k_{2}\,k_{3}+k_{3}^{2}\right)+k_{2}\,k_{3}^{2}\,(k_{2}+k_{3})^{2}\biggr{]}\,\eta_{\mathrm{e}}^{2}$
$\quad+i\,k_{1}\,k_{3}\,(k_{1}+k_{2}+k_{3})^{3}\,\left[k_{1}^{2}\,(3\,k_{2}+k_{3})+k_{1}\,\left(3\,k_{2}^{2}+11\,k_{2}\,k_{3}+k_{3}^{2}\right)+3\,k_{2}\,k_{3}\,(k_{2}+k_{3})\right]\,\eta_{\mathrm{e}}^{3}$
$\quad+k_{1}^{2}\,k_{2}\,k_{3}^{2}\,(k_{1}+k_{2}+k_{3})^{4}\,\eta_{\mathrm{e}}^{4}\biggr{\}}\,\mathrm{e}^{i\,(k_{1}+k_{2}+k_{3})\,\eta_{\mathrm{e}}}-9\,(k_{1}+k_{2}+k_{3})^{5}\,\mathrm{Ei}[i\,(k_{1}+k_{2}+k_{3})\,\eta_{\mathrm{e}}]\biggr{)}$
$\quad+{\bm{k}}_{2}\leftrightarrow{\bm{k}}_{3},$ (83d)
${\mathcal{G}}_{5}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=-\frac{iH_{{}_{\mathrm{I}}}}{256\,M_{{}_{\mathrm{Pl}}}\,\sqrt{k1^{3}\,k_{2}^{5}\,k_{3}^{5}\,\epsilon_{1\mathrm{i}}}\,\eta_{\mathrm{e}}^{3}}\,\left(\frac{\eta_{\mathrm{i}}}{\eta_{\mathrm{e}}}\right)^{3}\,\left[k_{1}^{4}-2\,k_{1}^{2}\,(k_{2}^{2}+k_{3}^{2})+k_{2}^{4}+6\,k_{2}^{2}\,k_{3}^{2}+k_{3}^{4}\right]$
×($\quad\times\biggl{(}\biggl{\{}-24\,i-24\,(k_{1}+k_{2}+k_{3})\,\eta_{\mathrm{e}}-6\,i\,\left[k_{1}^{2}-4\,k_{1}\,(k_{2}+k_{3})-k_{2}^{2}-4\,k_{2}\,k_{3}-k_{3}^{2}\right]\,\eta_{\mathrm{e}}^{2}$
@@
×$\quad\times\mathrm{Ei}[i\,(k_{1}+k_{2}+k_{3})\,\eta_{\mathrm{e}}]\biggr{)},$ (83e)
${\mathcal{G}}_{6}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=\frac{7\,i\,H_{{}_{\mathrm{I}}}}{768\,M_{{}_{\mathrm{Pl}}}\,\sqrt{k_{1}^{3}\,k_{2}^{5}\,k_{3}^{5}\,\epsilon_{1\mathrm{i}}}\,\eta_{\mathrm{e}}^{3}}\left(\frac{\eta_{\mathrm{i}}}{\eta_{\mathrm{e}}}\right)^{3}\,\left[k_{1}^{4}-2\,k_{1}^{2}\,(k_{2}^{2}+k_{3}^{2})+k_{2}^{4}+6\,k_{2}^{2}\,k_{3}^{2}+k_{3}^{4}\right]$
@@
×$\quad\times\,\mathrm{Ei}[i\,(k_{1}+k_{2}+k_{3})\,\eta_{\mathrm{e}}]\biggr{)},$ (83f)
@@ B.2 Case of $p=4$
${\mathcal{G}}_{2}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=-\frac{i\,H_{{}_{\mathrm{I}}}}{8\,M_{{}_{\mathrm{Pl}}}\,\sqrt{k_{1}\,k_{2}\,k_{3}\,\epsilon_{1\mathrm{i}}}}\,\left(\frac{\eta_{\mathrm{i}}}{\eta_{\mathrm{e}}}\right)^{2}\,(k_{1}^{2}-k_{2}^{2}-k_{3}^{2})\,\biggl{\{}\biggl{[}1-i\,(k_{1}+k_{2}+k_{3})\,\eta_{\mathrm{e}}$
$\quad-\frac{2\,k_{1}\,k_{2}\,k_{3}\,\eta_{\mathrm{e}}^{2}}{(k_{1}+k_{2}+k_{3})}\biggr{]}\,\mathrm{e}^{i\,(k_{1}+k_{2}+k_{3})\,\eta_{\mathrm{e}}}-(k_{1}^{2}+k_{2}^{2}+k_{3}^{2})\,\eta_{\mathrm{e}}^{2}\;\mathrm{Ei}[i\,(k_{1}+k_{2}+k_{3})\,\eta_{\mathrm{e}}]\biggr{\}},$ (84b)
${\mathcal{G}}_{3}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=\frac{i\,H_{{}_{\mathrm{I}}}}{32\,M_{{}_{\mathrm{Pl}}}\,\sqrt{k_{1}\,k_{2}\,k_{3}^{5}\,\epsilon_{1\mathrm{i}}}}\,\left(\frac{\eta_{\mathrm{i}}}{\eta_{\mathrm{e}}}\right)^{2}\,\left[2\,k_{3}^{2}\,(k_{1}^{2}+k_{2}^{2})+(k_{1}^{2}-k_{2}^{2})^{2}-3\,k_{3}^{4}\right]$
$\quad+(3\,k_{1}^{2}-3\,k_{2}^{2}-k_{3}^{2})\,\eta_{\mathrm{e}}^{2}\;\mathrm{Ei}[i\,(k_{1}+k_{2}+k_{3})\,\eta_{\mathrm{e}}]\bigg{\}}+{\bm{k}}_{2}\leftrightarrow{\bm{k}}_{3},$ (84c)
${\mathcal{G}}_{4}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=-\frac{H_{{}_{\mathrm{I}}}\,\sqrt{\epsilon_{1\mathrm{i}}}}{16\,M_{{}_{\mathrm{Pl}}}\,\sqrt{k_{1}^{5}\,k_{2}\,k_{3}^{5}}\,(k_{1}+k_{2}+k_{3})^{4}\,\eta_{\mathrm{i}}^{2}}\,\left[(k_{1}^{2}-k_{2}^{2})^{2}+2\,(k_{1}^{2}+k_{2}^{2})\,k_{3}^{2}-3\,k_{3}^{4}\right]$
×($\quad\times\biggl{(}\biggl{\{}-i\,\biggl{[}3\,k_{3}^{2}\,\left(9\,k_{1}^{2}+23\,k_{1}\,k_{2}+9\,k_{2}^{2}\right)+3\,k_{3}\,(k_{1}+k_{2})\,\left(5\,k_{1}^{2}+14\,k_{1}\,k_{2}+5\,k_{2}^{2}\right)$
$\quad+3\,(k_{1}+k_{2})^{2}\,\left(k_{1}^{2}+3\,k_{1}\,k_{2}+k_{2}^{2}\right)+19\,k_{3}^{3}\,(k_{1}+k_{2})+4\,k_{3}^{4}\biggr{]}$
$\quad-(k_{1}+k_{2}+k_{3})\,\biggl{[}3\,k_{3}^{2}\,\left(3\,k_{1}^{2}+11\,k_{1}\,k_{2}+3\,k_{2}^{2}\right)+3\,k_{3}\,(k_{1}+k_{2})\,\left(k_{1}^{2}+6\,k_{1}\,k_{2}+k_{2}^{2}\right)$
$\quad+7\,k_{3}^{3}\,(k_{1}+k_{2})+3\,k_{1}\,k_{2}\,(k_{1}+k_{2})^{2}+k_{3}^{4}\biggr{]}\,\eta_{\mathrm{e}}+i\,k_{3}\,(k_{1}+k_{2}+k_{3})^{2}\,\biggl{[}k_{3}\,\left(k_{1}^{2}+8\,k_{1}\,k_{2}+k_{2}^{2}\right)$
$\quad+k_{3}^{2}\,(k_{1}+k_{2})+3\,k_{1}\,k_{2}\,(k_{1}+k_{2})\biggr{]}\,\eta_{\mathrm{e}}^{2}+k_{1}\,k_{2}\,k_{3}^{2}\,(k_{1}+k_{2}+k_{3})^{3}\,\eta_{\mathrm{e}}^{3}\biggr{\}}\,\mathrm{e}^{i\,(k_{1}+k_{2}+k_{3})\,\eta_{\mathrm{e}}}$
$\quad+3\,i\,(k_{1}+k_{2}+k_{3})^{4}\,\textrm{Ei}[i\,(k_{1}+k_{2}+k_{3})\,\eta_{\mathrm{e}}]\biggr{)}+{\bm{k}}_{2}\leftrightarrow{\bm{k}}_{3},$ (84d)
${\mathcal{G}}_{5}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=-\frac{i\,H_{{}_{\mathrm{I}}}}{128\,M_{{}_{\mathrm{Pl}}}\,\sqrt{k_{1}\,k_{2}^{5}\,k_{3}^{5}\,\epsilon_{1\mathrm{i}}}\,\eta_{\mathrm{e}}^{2}}\,\left(\frac{\eta_{\mathrm{i}}}{\eta_{\mathrm{e}}}\right)^{2}\,\left[k_{1}^{4}-2\,k_{1}^{2}\,(k_{2}^{2}+k_{3}^{2})+k_{2}^{4}+6\,k_{2}^{2}\,k_{3}^{2}+k_{3}^{4}\right]$
@@
$\quad+\left[2\,k_{3}^{2}\,(k_{2}^{2}-3k_{1}^{2})+3\,(k_{1}^{2}-k_{2}^{2})^{2}+3\,k_{3}^{4}\right]\,\eta_{\mathrm{e}}^{4}\;\mathrm{Ei}[i\,(k_{1}+k_{2}+k_{3})\,\eta_{\mathrm{e}}]\biggr{)},$ (84e)
${\mathcal{G}}_{6}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=\frac{3\,i\,H_{{}_{\mathrm{I}}}}{128\,M_{{}_{\mathrm{Pl}}}\,\sqrt{k_{1}\,k_{2}^{5}\,k_{3}^{5}\,\epsilon_{1\mathrm{i}}}\,\eta_{\mathrm{e}}^{2}}\,\left(\frac{\eta_{\mathrm{i}}}{\eta_{\mathrm{e}}}\right)^{2}\,\left[k_{1}^{4}-2\,k_{1}^{2}\,(k_{2}^{2}+k_{3}^{2})+k_{2}^{4}+6\,k_{2}^{2}\,k_{3}^{2}+k_{3}^{4}\right]$